### v1

@@ 1 Introduction
For LRA, the Fourier-Motzkin elimination (FME) algorithm is the first quantifier elimination procedure, originally proposed by Fourier [fourier1826-fme] and rediscovered by Motzkin [motzkin1936-fme]. This algorithm can be understood as reducing a system of linear inequalities by removing variables one by one. Its importance primarily stems from its geometric interpretation, which corresponds to the projection of a polyhedron (described by linear inequalities) onto lower-dimensional subspaces [schrijver1998book-lp, Sect. 12.2]. However, a significant challenge of applying the FME algorithm is its high worst-case complexity, which is doubly exponential in the number of quantified variables. This high complexity results from the rapid, often redundant, proliferation of new constraints during the elimination steps [imbert90, imbert93ppcp, jing20casc]. For eliminating a block of existential quantifiers, recent works [promies24fm, promies25lmcs] provide a divide-and-conquer approach to improve the complexity to single exponential.
For NRA, the first quantifier elimination procedure was developed around the 1930s by Tarski in his seminal work [tarski1930]. While Tarski’s procedure is not elementary recursive, the first elementary recursive quantifier elimination procedure was developed by Collins in 1975 [collins75-cad], known as cylindrical algebraic decomposition (CAD). The CAD algorithm marks a milestone in real algebraic geometry and a comprehensive survey can be found in [caviness1998book-qe]. Ever since its origin, the CAD algorithm, as well as its various variants [mccallum88jsc, hong90issac, collins91, brown01jsc, chen09issac, hong12jsc, brown15issac, han16jsc, strzebonski16jsc, bradford16jsc, han17jsc], has remained the most authoritative procedure for NRA. It has found broad application in critical domains, including formal verification [kapur06jssc, rodriguez04issac, liu11emsoft, gan18tac, AnZLZY18], control synthesis [dorato97jsc, jirstrand97jsc, hong97jsc], and hybrid system analysis [anai01hscc, sturm11issac]. Similar to the FME algorithm, the CAD algorithm also suffers from a worst-case doubly exponential complexity, but in the number of all occurring variables [davenport88jsc]. Besides the complexity of internal algebraic operations, it is well known that the order of variables to be eliminated, called variable elimination ordering, has a huge impact on the practical performance when eliminating a block of existential quantifiers. Therefore, many heuristics based on sophisticated structural analysis [brown04issac, dolzmann04issac, bradford13cicm, li23jsc] and machine learning techniques have been applied to this task [huang14cicm, england19cicm, chen20icms, jia23nips].
@@
The study of parameterized algorithms offers a fine-grained analytical framework for tackling computationally hard problems by introducing a secondary measurement, $k$, known as the parameter [downey12book-parameterized, cygan15book-parameterized]. For example, a problem is classified as fixed-parameter tractable (FPT) if it admits an algorithm with a time complexity of $O(f(k)\cdot n^{c})$, where $n$ is the input size, $f$ is a computable function depending solely on $k$, and $c$ is a constant independent of $n$. In general, the goal of parameterized complexity is to conduct a finer-grained analysis of complexity by isolating the combinatorial explosion to a specific parameter.
When the input is a graph, treewidth is one of the most important parameters to measure its inherent complexity. Intuitively, treewidth quantifies how “tree-like” a graph is: a smaller treewidth value indicates a greater sparsity in the graph structure. The main advantage of treewidth is that a bounded treewidth allows a vast collection of classical NP-hard problems to become fixed-parameter tractable [robertson86-minor2, bodlaender97mfcs, cygan15book-parameterized]. This tractability is fundamentally realized by employing dynamic programming (DP) techniques executed over a tree structure related to the graph, called a tree decomposition [bodlaender88icalp-dp]. For constraint satisfiability problems, treewidth measures the structural complexity of variable dependencies and has been extensively studied [szeider03dam, marx10tc, fichte20cp, dong21stoc, gu22open-sdp, mallach25acta-ilp]. However, the role of treewidth in the quantifier elimination of first-order logic theories is largely untouched, possibly due to the inherent high complexity of these procedures.
@@
    We prove that, when the treewidth is a constant, the worst-time complexity of FME and CAD can be exponentially improved from doubly exponential to single exponential in the number of eliminated variables and all occurring variables, respectively. The core idea is the utilization of a special tree decomposition structure, called balanced tree decompositions [chatterjee18toplas].
@@
Our starting point is a parameterized algorithmic perspective, a methodology that is orthogonal to most existing techniques. The most closely related work is [li23jsc], which exploits the chordal structure of the primal graph to guide the variable elimination ordering in CAD. Since chordal graphs and tree decompositions are two sides of one coin [diestel00book-graph, Sec. 12.3], our algorithm in Sect. 3 can be viewed as a dual version of theirs. Nevertheless, utilizing treewidth offers significant algorithmic advantages. Specifically, it enables a dynamic programming framework that is more comprehensible and, by drawing upon concepts from parameterized complexity theory, a more fine-grained analysis of the complexity issues involved.
@@ 2 Preliminaries
Let $\mathbb{R}$ and $\mathbb{N}$ denote the set of real numbers and natural numbers, respectively. A vector of $n$ real variables is denoted by $(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$, also written as $\mathbf{x}$ for short. We denote by $\langle x_{\sigma(1)},\dots,x_{\sigma(n)}\rangle$ a linear ordering among these variables, where $\sigma$ is from the permutation group of size $n$. For example, $\langle x_{1},\dots,x_{n}\rangle$ denotes the natural order from $x_{1}$ to $x_{n}$. We assume readers are familiar with first-order logic and recommend the book [bradley07book-computation] for reference. For a set $C$ of atomic formulas, we denote by $\bigwedge C$ the conjunction of all atomic formulas in $C$.
@@ 2.1 LRA, NRA, and Quantifier Elimination
We denote the output by $\textsf{FME}(C,x_{r})$, which is a set of atomic formulas without $x_{r}$. If $C$ is taken to be the set of atomic formulas of an LRA formula $\Phi$ of the form Eq. 1, then $\Phi$ is equivalent to the following quantifier-free formula $\bigwedge\textsf{FME}(C,\langle x_{1},\dots,x_{m}\rangle)$, where $\textsf{FME}(C,\langle x_{1},\dots,x_{m}\rangle)$ denotes the output of recursively applying FME procedure on $C$ w.r.t. the ordering from $x_{1}$ to $x_{m}$. For a more detailed description of the algorithm, please refer to [daniel08book, Sec. 5.4].
@@
through the iterative application of a projection operator Proj. Each subsequent set $P_{i+1}=\textsf{Proj}(P_{i},x_{i})$ is defined in a lower-dimensional space by removing the variable $x_{i}$ and carries enough information about $P_{i}$. There exist several different definitions of the projection operator, and in our analysis, we use McCallum’s projection operator [mccallum88jsc]. After the projection phase, the lifting phase then constructs the sign-invariant cells from 1-dimension to higher dimensions. We provide a more detailed description of the CAD algorithm [arnon84jsc, jirstrand95, basu06book-rag] in Sect. 0.A.1, which may not affect the understanding of this paper.
@@ 2.2 Treewidth and Tree Decomposition
In this part, we introduce basic concepts related to the sparsity of graphs. We recommend referring to [cygan15book-parameterized, Ch. 7] for an in-depth investigation of this topic.
@@
Consider an LRA formula $\Phi$ of the form Eq. 1 and contains existentially quantified variables $(x_{1},\dots,x_{8})$. The portion of $\Phi$ related to $x_{1}$ is displayed in Sect. 2.2; the full formula containing 20 atomic formulas is in Sect. 0.A.2. The atomic formulas in $\Phi$ display a certain “sparsity” pattern—each variable only occurs in a small portion of atomic formulas. For example, variable $x_{1}$ only occurs in three atomic formulas, namely $x_{1}+2x_{2}+3x_{3}\leq 20$, $x_{1}-x_{2}+2x_{3}\geq-5$, and $x_{1}-4x_{2}\leq 0$. The primal graph of $\Phi$ is shown in Fig. 2.
@@
Tree Decomposition and Treewidth [halin76-sfunction, robertson84-minor3, robertson86-minor2]. Given an undirected graph $G=(V,E)$, a tree decomposition of $G$ is a tree $T=(\mathcal{B},E_{T})$ satisfying the following conditions: (1) Each node $b\in\mathcal{B}$ of $T$, also called a bag, is associated to a subset of the vertex set $V_{b}\subseteq V$; (2) The bags cover the entire vertex set $V$ of $G$, i.e., $\bigcup_{b\in\mathcal{B}}V_{b}=V$; (3) For every edge $e=(u,v)\in E$, there exists a bag $b$ such that both $u\in V_{b}$ and $v\in V_{b}$; and (4) For every three bags $b_{1},b_{2},b_{3}\in\mathcal{B}$, if $b_{3}$ is on the path from $b_{1}$ to $b_{2}$ in $T$, then $V_{b_{1}}\cap V_{b_{2}}\subseteq V_{b_{3}}$ (equivalently, every vertex $v\in V$ appears in a connected subtree of $T$).
@@
The width of a tree decomposition $T$, denoted by $w(T)$, is defined to be the size of the largest bag minus one, i.e., $w(T)=\max_{b\in\mathcal{B}}|V_{b}|-1$. The treewidth of the graph $G$, denoted by $tw(G)$, is defined to be the minimum width among all possible tree decompositions of $G$. A tree decomposition with width $tw(G)$ is called optimal. For example, Fig. 3 presents an optimal tree decomposition of the primal graph in Exmp. 1 of width $2$. The numbers in the labels refer to the variable indexes, for instance, the node labeled $\{1,2,3\}$ in Fig. 3 represents the bag $\{x_{1},x_{2},x_{3}\}$. For a primal graph $G_{\Phi}$, since every clique is contained in at least one bag in its tree decomposition [cygan15book-parameterized], the largest number of quantified variables in an atomic formula gives a lower bound for $tw(G_{\Phi})$.
@@
A smaller treewidth indicates a greater sparsity. In general, it is NP-complete to determine whether a given graph $G$ has treewidth at most a given variable $k$ [stefan87jadm]. However, Bodlaender showed that when $k$ is fixed, there exists a linear-time algorithm to test whether $G$ has treewidth at most $k$, and if so, to construct an optimal tree-decomposition with $O(n)$ nodes [bodlaender96joc]. In this paper, we rely on an external algorithm [tamaki19jco-td] to compute tree decompositions. There are also many other well-optimized tools, see [dell17ipec-pace] for a survey.
@@
Nice Tree Decomposition [kloks94book-treewidth]. A tree decomposition $T=(\mathcal{B},E_{T})$ of a graph $G$ is nice if $T$ is rooted at a bag $r\in\mathcal{B}$ with $V_{r}=\emptyset$ and each bag of $T$ is of one of the following four types: (1) leaf bag is a bag $b$ having no children and $|V_{b}|=1$; (2) A join bag is a bag $b$ having exactly two children $b_{1}$ and $b_{2}$ such that $V_{b}=V_{b_{1}}=V_{b_{2}}$; (3) An introduce bag is a bag $b$ having exactly one child $b^{\prime}$ such that $|V_{b}|=|V_{b^{\prime}}|+1$ and $V_{b^{\prime}}\subseteq V_{b}$; (4) A forget bag is a bag $b$ having exactly one child $b^{\prime}$ such that $|V_{b}|=|V_{b^{\prime}}|-1$ and $V_{b}\subseteq V_{b^{\prime}}$. As an example, Fig. 4 gives a nice tree decomposition based on the tree decomposition in Fig. 3.
@@
The concept of nice tree decomposition is introduced as a canonical form for representing tree decompositions. When the treewidth is a constant, a tree decomposition can be transformed into a nice tree decomposition in linear time, with the same width but a linear increase in size [cygan15book-parameterized, Lem. 7.4], which helps design dynamic programming algorithms and conduct computational analysis.
@@ 3.1 A Dynamic Programming Framework for FME
Using tree decompositions, we can re-explain the process of eliminating $x_{1}$ in a tree-like manner. Consider the nice tree decomposition presented in Fig. 4, the process of eliminating $x_{1}$ consists of two steps. In the first step, we assign each atomic formula of $\Phi$ to a bag in the nice tree decomposition depending on the involved variables. This is described by the initialization of a function $\mathcal{I}$ that maps each bag to a set of atomic formulas. For example, for the three bags located at the left-most branch of the nice tree decomposition in Fig. 4, we set $\mathcal{I}(\{x_{1}\})\leftarrow\{\}$ (meaning there is no atomic formulas containing only $x_{1}$), $\mathcal{I}(\{x_{1},x_{2}\})\leftarrow\{x_{1}-4x_{2}\leq 0\}$, and $\mathcal{I}(\{x_{1},x_{2},x_{3}\})\leftarrow\{x_{1}-x_{2}+2x_{3}\geq-5,x_{1}+2x_{2}+3x_{3}\leq 20\}$. Note that the atomic formula $x_{1}-4x_{2}\leq 0$ can also be assigned to $\mathcal{I}(\{x_{1},x_{2},x_{3}\})$ without affecting the result.
@@
3 Construct the primal graph $G_{\Phi}=(V_{\Phi},E_{\Phi})$ of $\Phi$;
4 Compute a nice tree decomposition $T=(\mathcal{B},E_{T})$ of $G_{\Phi}$;
5 $\triangleright$ Initialize $\mathcal{I}$ from top down.
6 $C\leftarrow$ atomic formulas in $\Phi$;
7 Initialize an empty queue of bags $q=\langle\rangle$;
8 $q.\texttt{enqueue}(r_{T})$ ;
9 // add to the end of the queue
10 while $q$ is not empty do
11 $b\leftarrow q.\texttt{dequeue}()$ ;
12 // remove from the front of the queue
13 $q.\texttt{enqueue}(b.\texttt{children})$;
14 // break ties with any heuristics
15 if $b.\textnormal{{type}}=\textsf{leaf}$ then
16    $\mathcal{I}(b)\leftarrow\{\}$ ;
17    // if $b$ is a leaf bag
18
19 else
20    $\mathcal{I}(b)\leftarrow\{\varphi\in C\mid\texttt{var}(\varphi)\subseteq b\}$ ;
21    // if $b$ is of other types
22    $C\leftarrow C\setminus\mathcal{I}(b)$;
23
24   end if
25
26 end while
[…]
28 $q.\texttt{enqueue}(\mathcal{B}.\texttt{leaves})$ ;
29 while $q$ is not empty do
30 $b\leftarrow q.\texttt{dequeue}()$;
31 $q.\texttt{enqueue}(b.\texttt{parent})$;
32   Compute $\mathcal{V}(b)$ according to Eq. 3;
33
34 end while

### v2

@@ 1 Introduction
For LRA, the Fourier-Motzkin elimination (FME) algorithm is the first quantifier elimination procedure, originally proposed by Fourier [33] and rediscovered by Motzkin [61]. This algorithm can be understood as reducing a system of linear inequalities by removing variables one by one. Its importance primarily stems from its geometric interpretation, which corresponds to the projection of a polyhedron (described by linear inequalities) onto lower-dimensional subspaces [68, Sect. 12.2]. However, a significant challenge of applying the FME algorithm is its high worst-case complexity, which is doubly exponential in the number of quantified variables. This high complexity results from the rapid, often redundant, proliferation of new constraints during the elimination steps [43, 44, 51, 47]. For eliminating a block of existential quantifiers, recent works [62, 63] provide a divide-and-conquer approach to improve the complexity to single exponential.
For NRA, the first quantifier elimination procedure was developed around the 1930s by Tarski in his seminal work [73]. While Tarski’s procedure is not elementary recursive, the first elementary recursive quantifier elimination procedure was developed by Collins in 1975 [21], known as cylindrical algebraic decomposition (CAD). The CAD algorithm marks a milestone in real algebraic geometry and a comprehensive survey can be found in [16]. Ever since its origin, the CAD algorithm, as well as its various variants [60, 41, 20, 13, 18, 39, 15, 38, 69, 10, 37], has remained the most authoritative procedure for NRA. It has found broad application in critical domains, including formal verification [50, 66, 55, 34, 1], control synthesis [28, 49, 40], and hybrid system analysis [2, 70]. Similar to the FME algorithm, the CAD algorithm also suffers from a worst-case doubly exponential complexity, but in the number of all occurring variables [23]. Besides the complexity of internal algebraic operations, it is well known that the order of variables to be eliminated, called variable elimination ordering, has a huge impact on the practical performance when eliminating a block of existential quantifiers. Therefore, many heuristics based on sophisticated structural analysis [14, 26, 11, 54] and machine learning techniques have been applied to this task [42, 31, 19, 46].
@@
The study of parameterized algorithms offers a fine-grained analytical framework for tackling computationally hard problems by introducing a secondary measurement, $k$, known as the parameter [29, 22]. For example, a problem is classified as fixed-parameter tractable (FPT) if it admits an algorithm with a time complexity of $O(f(k)\cdot n^{c})$, where $n$ is the input size, $f$ is a computable function depending solely on $k$, and $c$ is a constant independent of $n$. In general, the goal of parameterized complexity is to conduct a finer-grained analysis of complexity by isolating the combinatorial explosion to a specific parameter.
When the input is a graph, treewidth is one of the most important parameters to measure its inherent complexity. Intuitively, treewidth quantifies how “tree-like” a graph is: a smaller treewidth value indicates a greater sparsity in the graph structure. The main advantage of treewidth is that a bounded treewidth allows a vast collection of classical NP-hard problems to become fixed-parameter tractable [64, 9, 22]. This tractability is fundamentally realized by employing dynamic programming (DP) techniques executed over a tree structure related to the graph, called a tree decomposition [7]. For constraint satisfiability problems, treewidth measures the structural complexity of variable dependencies and has been extensively studied [71, 58, 32, 27, 35, 56]. However, the role of treewidth in the quantifier elimination of first-order logic theories is largely untouched, possibly due to the inherent high complexity of these procedures.
@@
    We prove that, when the treewidth is a constant, the worst-time complexity of FME and CAD can be exponentially improved from doubly exponential to single exponential in the number of eliminated variables and all occurring variables, respectively. The core idea is the utilization of a special tree decomposition structure, called balanced tree decompositions [17].
@@
Our starting point is a parameterized algorithmic perspective, a methodology that is orthogonal to most existing techniques. The most closely related work is [54], which exploits the chordal structure of the primal graph to guide the variable elimination ordering in CAD. Since chordal graphs and tree decompositions are two sides of one coin [25, Sec. 12.3], our algorithm in Sect. 3 can be viewed as a dual version of theirs. Nevertheless, utilizing treewidth offers significant algorithmic advantages. Specifically, it enables a dynamic programming framework that is more comprehensible and, by drawing upon concepts from parameterized complexity theory, a more fine-grained analysis of the complexity issues involved.
@@ 2 Preliminaries
Let $\mathbb{R}$ and $\mathbb{N}$ denote the set of real numbers and natural numbers, respectively. A vector of $n$ real variables is denoted by $(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$, also written as $\mathbf{x}$ for short. We denote by $\langle x_{\sigma(1)},\dots,x_{\sigma(n)}\rangle$ a linear ordering among these variables, where $\sigma$ is from the permutation group of size $n$. For example, $\langle x_{1},\dots,x_{n}\rangle$ denotes the natural order from $x_{1}$ to $x_{n}$. We assume readers are familiar with first-order logic and recommend the book [12] for reference. For a set $C$ of atomic formulas, we denote by $\bigwedge C$ the conjunction of all atomic formulas in $C$.
@@ 2.1 LRA, NRA, and Quantifier Elimination
We denote the output by $\textsf{FME}(C,x_{r})$, which is a set of atomic formulas without $x_{r}$. If $C$ is taken to be the set of atomic formulas of an LRA formula $\Phi$ of the form Eq. 1, then $\Phi$ is equivalent to the following quantifier-free formula $\bigwedge\textsf{FME}(C,\langle x_{1},\dots,x_{m}\rangle)$, where $\textsf{FME}(C,\langle x_{1},\dots,x_{m}\rangle)$ denotes the output of recursively applying FME procedure on $C$ w.r.t. the ordering from $x_{1}$ to $x_{m}$. For a more detailed description of the algorithm, please refer to [53, Sec. 5.4].
@@
through the iterative application of a projection operator Proj. Each subsequent set $P_{i+1}=\textsf{Proj}(P_{i},x_{i})$ is defined in a lower-dimensional space by removing the variable $x_{i}$ and carries enough information about $P_{i}$. There exist several different definitions of the projection operator, and in our analysis, we use McCallum’s projection operator [60]. After the projection phase, the lifting phase then constructs the sign-invariant cells from 1-dimension to higher dimensions. We provide a more detailed description of the CAD algorithm [4, 48, 67] in Sect. 0.A.1. which may not affect the understanding of this paper.
@@ 2.2 Treewidth and Tree Decomposition
In this part, we introduce basic concepts related to the sparsity of graphs. We recommend referring to [22, Ch. 7] for an in-depth investigation of this topic.
@@
Consider an LRA formula $\Phi$ of the form Eq. 1 and contains existentially quantified variables $(x_{1},\dots,x_{8})$. The portion of $\Phi$ related to $x_{1}$ is displayed in Sect. 2.2; the complete formula containing 20 atomic formulas is in the full version Sect. 0.A.2. The atomic formulas in $\Phi$ display a certain “sparsity” pattern—each variable only occurs in a small portion of atomic formulas. For example, variable $x_{1}$ only occurs in three atomic formulas, namely $x_{1}+2x_{2}+3x_{3}\leq 20$, $x_{1}-x_{2}+2x_{3}\geq-5$, and $x_{1}-4x_{2}\leq 0$. The primal graph of $\Phi$ is shown in Fig. 2.
@@
Tree Decomposition and Treewidth [36, 65, 64]. Given an undirected graph $G=(V,E)$, a tree decomposition of $G$ is a tree $T=(\mathcal{B},E_{T})$ satisfying the following conditions: (1) Each node $b\in\mathcal{B}$ of $T$, also called a bag, is associated to a subset of the vertex set $V_{b}\subseteq V$; (2) The bags cover the entire vertex set $V$ of $G$, i.e., $\bigcup_{b\in\mathcal{B}}V_{b}=V$; (3) For every edge $e=(u,v)\in E$, there exists a bag $b$ such that both $u\in V_{b}$ and $v\in V_{b}$; and (4) For every three bags $b_{1},b_{2},b_{3}\in\mathcal{B}$, if $b_{3}$ is on the path from $b_{1}$ to $b_{2}$ in $T$, then $V_{b_{1}}\cap V_{b_{2}}\subseteq V_{b_{3}}$ (equivalently, every vertex $v\in V$ appears in a connected subtree of $T$).
@@
The width of a tree decomposition $T$, denoted by $w(T)$, is defined to be the size of the largest bag minus one, i.e., $w(T)=\max_{b\in\mathcal{B}}|V_{b}|-1$. The treewidth of the graph $G$, denoted by $tw(G)$, is defined to be the minimum width among all possible tree decompositions of $G$. A tree decomposition with width $tw(G)$ is called optimal. For example, Fig. 3 presents an optimal tree decomposition of the primal graph in Exmp. 1 of width $2$. The numbers in the labels refer to the variable indexes, for instance, the node labeled $\{1,2,3\}$ in Fig. 3 represents the bag $\{x_{1},x_{2},x_{3}\}$. For a primal graph $G_{\Phi}$, since every clique is contained in at least one bag in its tree decomposition [22], the largest number of quantified variables in an atomic formula gives a lower bound for $tw(G_{\Phi})$.
@@
A smaller treewidth indicates a greater sparsity. In general, it is NP-complete to determine whether a given graph $G$ has treewidth at most a given variable $k$ [3]. However, Bodlaender showed that when $k$ is fixed, there exists a linear-time algorithm to test whether $G$ has treewidth at most $k$, and if so, to construct an optimal tree-decomposition with $O(n)$ nodes [8]. In this paper, we rely on an external algorithm [72] to compute tree decompositions. There are also many other well-optimized tools, see [24] for a survey.
@@
Nice Tree Decomposition [52]. A tree decomposition $T=(\mathcal{B},E_{T})$ of a graph $G$ is nice if $T$ is rooted at a bag $r\in\mathcal{B}$ with $V_{r}=\emptyset$ and each bag of $T$ is of one of the following four types: (1) leaf bag is a bag $b$ having no children and $|V_{b}|=1$; (2) A join bag is a bag $b$ having exactly two children $b_{1}$ and $b_{2}$ such that $V_{b}=V_{b_{1}}=V_{b_{2}}$; (3) An introduce bag is a bag $b$ having exactly one child $b^{\prime}$ such that $|V_{b}|=|V_{b^{\prime}}|+1$ and $V_{b^{\prime}}\subseteq V_{b}$; (4) A forget bag is a bag $b$ having exactly one child $b^{\prime}$ such that $|V_{b}|=|V_{b^{\prime}}|-1$ and $V_{b}\subseteq V_{b^{\prime}}$. As an example, Fig. 4 gives a nice tree decomposition based on the tree decomposition in Fig. 3.
@@
The concept of nice tree decomposition is introduced as a canonical form for representing tree decompositions. When the treewidth is a constant, a tree decomposition can be transformed into a nice tree decomposition in linear time, with the same width but a linear increase in size [22, Lem. 7.4], which helps design dynamic programming algorithms and conduct computational analysis.
@@ 3.1 A Dynamic Programming Framework for FME
Using tree decompositions, we can re-explain the process of eliminating $x_{1}$ in a tree-like manner. Consider the nice tree decomposition presented in Fig. 4, the process of eliminating $x_{1}$ consists of two steps. In the first step, we assign each atomic formula of $\Phi$ to a bag in the nice tree decomposition depending on the involved variables. This is described by the initialization of a function $\mathcal{I}$ that maps each bag to a set of atomic formulas. For example, for the three bags located at the left-most branch of the nice tree decomposition in Fig. 4, we set $\mathcal{I}(\{x_{1}\})\leftarrow\{\}$ (meaning there are no atomic formulas containing only $x_{1}$), $\mathcal{I}(\{x_{1},x_{2}\})\leftarrow\{x_{1}-4x_{2}\leq 0\}$, and $\mathcal{I}(\{x_{1},x_{2},x_{3}\})\leftarrow\{x_{1}-x_{2}+2x_{3}\geq-5,x_{1}+2x_{2}+3x_{3}\leq 20\}$. Note that the atomic formula $x_{1}-4x_{2}\leq 0$ can also be assigned to $\mathcal{I}(\{x_{1},x_{2},x_{3}\})$ without affecting the result.
@@
Construct the primal graph $G_{\Phi}=(V_{\Phi},E_{\Phi})$ of $\Phi$;
Compute a nice tree decomposition $T=(\mathcal{B},E_{T})$ of $G_{\Phi}$;
$\triangleright$ Initialize $\mathcal{I}$ from top down.
$C\leftarrow$ atomic formulas in $\Phi$;
Initialize an empty queue of bags $q=\langle\rangle$;
$q.\texttt{enqueue}(r_{T})$ ;
// add to the end of the queue
while $q$ is not empty do
 $b\leftarrow q.\texttt{dequeue}()$ ;
 // remove from the front of the queue
 $q.\texttt{enqueue}(b.\texttt{children})$;
 // break ties with any heuristics
 if $b.\textnormal{{type}}=\textsf{leaf}$ then
    $\mathcal{I}(b)\leftarrow\{\}$ ;
    // if $b$ is a leaf bag
 else
    $\mathcal{I}(b)\leftarrow\{\varphi\in C\mid\texttt{var}(\varphi)\subseteq b\}$ ;
    // if $b$ is of other types
    $C\leftarrow C\setminus\mathcal{I}(b)$;
   end if
end while
$\triangleright$ Recursively compute $\mathcal{V}$ from bottom up.
$q.\texttt{enqueue}(\mathcal{B}.\texttt{leaves})$ ;
while $q$ is not empty do
 $b\leftarrow q.\texttt{dequeue}()$;
 $q.\texttt{enqueue}(b.\texttt{parent})$;
   Compute $\mathcal{V}(b)$ according to Eq. 3;
end while